A car starts from rest and attains the velocity of 60km/h in 5 minutes. It travels at the constant velocity for 1 minute and then decreases its speed to 30 km/h in another 5 minutes. Calculate the acceleration in si units in both the cases.
1
Expert's answer
2020-06-03T12:06:00-0400

The acceleration


"a=\frac{\Delta v}{\Delta t}"

"a_1=\frac{(60-0)\times \frac{1000\:\rm m}{3600\:\rm s}}{5\times 60\:\rm s}=0.056\:\rm m/s^2"

"a_2=\frac{(30-60)\times \frac{1000\:\rm m}{3600\:\rm s}}{5\times 60\:\rm s}=-0.028\:\rm m/s^2"